OpenStudy (anonymous):

Math Analysis: Prove the given identity cot 0 - tan 0/sin 0*cos 0=csc^2 0 - sec^2 0

OpenStudy (anonymous):

cot 0 - tan 0/sin 0*cos 0 is a bit ambiguous... do you mean \[\cot \theta-\frac{\tan \theta}{\sin \theta * \cos \theta}\]

OpenStudy (anonymous):

yah but cot 0 and tan 0 share common denominators with each other :P

OpenStudy (anonymous):

or this... \cot \theta-\frac{\tan \theta}{\sin \theta}*cos theta

OpenStudy (anonymous):